OpenStudy (anonymous):

Cheyenne drove 3,000 miles in 75 hours. At this rate, how long will I take her to drive 4,000 miles?

OpenStudy (danjs):

\[\frac{ 3000 mi}{ 75 hr } = \frac{ 4000 mi }{ x hr? }\]

OpenStudy (danjs):

\[x = \frac{ 4000 * 75 }{ 3000 }\]

OpenStudy (anonymous):

100

OpenStudy (anonymous):

100 hours using proprtions